(a) The random variable Y has a Poisson distribution and is such that P(Y = 0) = P(Y = 1). What is P(Y 2 = 1)?

(b) Cars arrive at a toll both according to a Poisson process with mean 80 cars per hour. If the attendant makes a one-minute phone call, what is the probability that at least 1 car arrives during the call?


1
Expert's answer
2022-04-13T12:10:45-0400

a) Poisson distribution is described by the formula "P(y)=\lambda^y e^{-\lambda}/y!" ,

where "\lambda" is distribution's mean

"P(0)=P(1)\Lrarr \lambda^0e^{-\lambda}/0!=\lambda^1e^{-\lambda}/1!\Lrarr\lambda=1\\
P(0)=P(1)=1/e,\space P(2)=1/(2!\cdot e)=1/(2e)"


b) Mean for 1-minute interval is 80 / 60 = 4 / 3.

Probability that neither car will arrive:

"P(0)=((4\/3)^0e^{-4\/3})\/0!=e^{-4\/3}"

Probability that at least one car will arrive is equal to "1-P(0)=1-e^{-4/3}\approx0.7364=73.64\%"